# Recovering contact forms from boundary data

### v5

@@ 6. Non-squeezing of $2n$-symplectic volumes, Morse wrinkles of Reeb fields, and holography
Using that $\beta_{1}=\Psi^{\ast}(\beta_{2})$, we see that $\Psi_{\ast}(v_{1})=v_{2}|_{\Psi(X_{1})}$. Moreover, the combinatorial tangency pattern of every $v_{1}$-trajectory $\gamma_{1}$ to $\partial_{1}X_{1}$ coincides with the tangency pattern of the portion $\Psi(\gamma_{1})$ of the unique $v_{2}$-trajectory $\gamma_{2}\supset\Psi(\gamma_{1})$ to $\Psi(\partial_{1}X)$. Note that the tangency pattern of $\gamma_{2}$ to $\Psi(\partial_{1}X)$ could be different.
@@
Now we are going to validate the equatorial inequality (see Fig. 1). Let $\mathcal{W}_{2}(\Psi)$ denote the $2n$-dimensional set comprised of $v_{2}$-trajectories through the locus $\Psi(\partial_{2}X_{1}(v_{1}))\coprod\partial_{2}X_{2}(v_{2})$. Consider the set (see Fig.  1)
@@
along which the continuous dependence of the $v_{2}$-flow on the initial conditions in $\Psi(\partial_{1}^{+}X_{1}(v_{1}))$ holds. Let $\mathcal{V}^{\circ}_{1,\kappa}(\Psi)$ be a connected component of the set $\mathcal{V}_{1}^{\circ}(\Psi)$ and let $\mathcal{V}_{1,\kappa}(\Psi)$ be the closure of $\mathcal{V}_{1,\kappa}^{\circ}(\Psi)$. We denote by $\partial^{+}_{1}X^{\circ}_{1,\kappa}(v_{1})\subset\partial_{1}^{+}X(v_{1})$ the $\Psi$-preimage of $\mathcal{V}^{\circ}_{1,\kappa}(\Psi)$.
@@
Let $\mathcal{Y}_{\kappa}(\Psi)=_{\mathsf{def}}\mathcal{X}_{\kappa}(\Psi)\cap\partial_{1}^{+}X_{2}(v_{2})$ be the set where the $v_{2}$-trajectories through $\mathcal{V}_{1,\kappa}(\Psi)$ terminate. Then the boundary of the set $\mathcal{X}_{\kappa}(\Psi)$ consists of three parts: $\mathcal{V}_{1,\kappa}(\Psi)),\,\mathcal{Y}_{\kappa}(\Psi)$, and the rest, denoted by $\mathcal{Z}_{\kappa}(\Psi)$. The $2n$-dimensional Lebesgue measures of mutual intersections of these three parts is zero.
@@
Adding the inequalities for all $\kappa$ and using that $\{\partial_{1}^{+}X^{\circ}_{1,\kappa}(v_{1})\}_{\kappa}$ form a partition of $\partial_{1}^{+}X_{1}(v_{1})$, modulo the set of $2n$-dimensional measure zero, we get
@@
[…]
where $\Gamma(\sim)$ is the Gamma function. In particular, for $X$ being a ball of radius $r$, we get the following constraints on the injectivity radius $r$ of a contact embedding:
$\frac{\pi^{n+\frac{1}{2}}}{\Gamma(n+\frac{3}{2})}\;r^{2n+1}\leq\int_{Y}\beta_{Y}\wedge(d\beta_{Y})^{n},$
$r\leq\mathsf{diam}_{\mathcal{R}}(\beta_{Y}),$
[…]
$\diamondsuit$
@@
Consider a solid $X$ shaped as a sand clock (see Fig. 2) which is vertically alined with the $z$-axis in the $xyz$-space $\mathbb{R}^{3}$. Let $\beta=dz+x\,dy$. We consider a $2$-parameter family of transformations $\{A_{\lambda\mu}:x\to\lambda x,y\to\lambda y,z\to\mu z\}_{\lambda,\mu\in\mathbb{R}_{+}}$. We compare two contact $3$-folds $(X,\beta)$ and $(A_{\lambda\mu}(X),\beta)$. Then $c^{\bullet}(\Psi,v_{2})\geq 2$.
@@
Let $v_{\beta}$ be a boundary generic Reeb vector field of a contact form $\beta$ on a compact $(2n+1)$-manifold $X$. For any boundary generic $v_{\beta}$, $\partial_{j}X(v_{\beta})=\partial(\partial_{j-1}^{+}X(v_{\beta}))$ for all $j\geq 1$. Thus, for an odd $j$, by Stokes’ theorem,
[…]
@@
$\kappa_{j}^{+}(\beta)$ $=_{\mathsf{def}}\;$ $\int_{\partial_{j}^{+}X(v_{\beta})}(d\beta)^{\frac{2n+1-j}{2}}\;\;\;(=\kappa_{j+1}(\beta)),$
[…]
@@
(6.6) $\kappa_{j}(\beta)$ $=_{\mathsf{def}}\;$ $\int_{\partial_{j}X(v)}\beta\wedge(d\beta)^{\frac{2n-j}{2}}\;\;\;(=\kappa^{+}_{j-1}(\beta))\text{\; for an even }j\in[2,2n],$
Of course, these quantities depend on the choice of the contact form $\beta$ that represents a given contact structure $\xi$ and whose Reeb vector field $v_{\beta}$ is boundary generic. They measure how “wrinkled” $\partial_{1}X$ is in relation to the Reeb flow $v_{\beta}$. By the very definition of $\kappa_{j}^{+}(\beta),\kappa_{j}(\beta)$, they belong to the category of “boundary data”.
Note that the inequality (6.2) can be written as $\kappa^{+}_{0}(\beta)\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\kappa^{+}_{1}(\beta)$, and the inequality (6.3) as $\kappa^{+}_{0}(\beta)\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\kappa_{2}(\beta)$, where $\kappa^{+}_{0}(\beta)=_{\mathsf{def}}\int_{X}\beta\wedge(d\beta)^{n}$.
@@ 6.1. Invariance of the volumes $\{\kappa_{j}(\beta)\}$ under special deformations of $\beta$
Let =$\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ be a family of contact forms on a compact connected smooth $(2n+1)$-manifold $X$ with boundary.
@@
for a family of smooth functions $\eta_{t}:\partial_{1}X\to\mathbb{R}$.
$\bullet$ We say the family $\mathcal{B}$ is variationally boundary $(2n+1-j)$-exact if, for some smooth family of $(2n-j)$-forms $\{\zeta_{t}\}_{t}$ on $\partial_{1}X$,
(6.7) $\Big{\{}\frac{\partial}{\partial t}\Big{[}\beta_{t}\wedge(d\beta_{t})^{\frac{2n-j}{2}}\Big{]}\Big{\}}\Big{|}_{\partial_{1}X}=d\zeta_{t},\text{ where }j\in[2,2n]\text{ is even, }$,
(6.8) $\Big{\{}\frac{\partial}{\partial t}\Big{[}(d\beta_{t})^{\frac{2n+1-j}{2}}\Big{]}\Big{\}}\Big{|}_{\partial_{1}X}=d\zeta_{t},\text{ where }j\in[1,2n-1]\text{ is odd }.$
@@
Any variationally boundary exact family $\mathcal{B}$ of contact forms is variationally boundary $(2n+1-j)$-exact.
@@
Note that, for any $t$-family differential forms, their exterior differential $d=_{\mathsf{def}}d_{t}$ commutes with the partial derivative operator $\frac{\partial}{\partial t}$ and with the restriction of forms to $\partial_{1}X$. Consider the derivative
@@
Since $\frac{\partial}{\partial t}\Big{[}(d\beta_{t})^{k}\Big{]}\Big{|}_{\partial_{1}X}=0$, this formula collapses to $\frac{\partial}{\partial t}\Big{[}\beta_{t}\Big{]}\wedge(d\beta_{t})^{k}\Big{|}_{\partial_{1}X}$.
Replacing $\frac{\partial}{\partial t}\Big{[}\beta_{t}\Big{]}\Big{|}_{\partial_{1}X}$ with $d\eta_{t}$ once more, we get
@@
Let =$\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ be a variationally boundary $(2n+1-j)$-exact family of contact forms on a compact connected smooth $(2n+1)$-manifold $X$. Assume that their Reeb vector fields $v_{\beta_{t}}$ are boundary generic for $t=0,1$, and for almost all $t\in(0,1)$.
$\bullet$ Then the integral $\kappa_{j}(\beta_{t})$ ($j\equiv 0\mod 2$) from (6.6) and $\kappa_{j-1}^{+}(\beta_{t})$ ($j\equiv 0\mod 2$) from (6) are $t$-independent. If, in addition, the property (6.9) holds, then $\kappa^{+}_{j}(\beta_{t})$ ($j\equiv 0\mod 2$) is also invariant within $\mathcal{B}$.
$\bullet$ If =$\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ is variationally boundary exact family, then $\{\kappa_{j}(\beta_{t})\}$ are $\mathcal{B}$-invariant for all even $j\in[2,2n]$. If in addition, (6.9) holds for the function $\eta_{t}$, then $\kappa^{+}_{j}(\beta_{t})$ ($j\equiv 0\mod 2$) is $\mathcal{B}$-invariant.
@@
$=\int_{\mathcal{Y}_{j}}(d_{t}\beta_{t})^{s+1}\;+\int_{\mathcal{Y}_{j}}\frac{\partial}{\partial t}\Big{[}\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big{]}\wedge dt$
Each of these two integrals vanishes. Indeed, $dt$ is not present in the $(2s+2)$-form $(d_{t}\beta_{t})^{s+1}$. Thus its restriction to the $(2s+2)$-chain $\mathcal{Y}_{j}\subset\partial_{1}X\times[0,1]$ vanishes. The second integral is also zero by the variationally boundary $(2n+1-j)$-exact hypothesis:
[…]
@@
Using Stokes’ theorem identity $\kappa_{j}(\beta)=\kappa_{j-1}^{+}(\beta)$ ($j\equiv 0\mod 2$), we conclude that, for odd $j\in[1,2n-1]$, the quantities $\{\kappa_{j}^{+}(\beta)\}_{j}$ are $\mathcal{B}$-deformation invariants as well.
If we add property (6.9) to property (6.7), then a similar computation shows that $\kappa^{+}_{j}(\beta_{1})=\kappa^{+}_{j}(\beta_{0})$ for an even $j$. This time, Stokes’ theorem is applied to $\mathcal{Y}_{j}^{+}$. Its boundary consists of tree parts $\partial_{j}^{+}X(v_{0}),\partial_{j}^{+}X(v_{1})$, and their complement $\delta\mathcal{Y}_{j}^{+}=\mathcal{Y}_{j+1}$. Therefore,
@@
$=\int_{\mathcal{Y}^{+}_{j}}(d_{t}\beta_{t})^{s+1}\;+\int_{\mathcal{Y}^{+}_{j}}\frac{\partial}{\partial t}\Big{[}\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big{]}\wedge dt\;\pm\;\int_{\delta\mathcal{Y}^{+}_{j}}\beta_{t}\wedge(d_{t}\beta_{t})^{s}.$
@@
[…]
@@
[…]
@@
[…]
@@
Then the integrals $\{\kappa_{j}^{+}(\beta_{X})\}_{j}$ from (6) and $\{\kappa_{j}(\beta_{X})\}_{j}$ from (6.6) can be recovered from the $\Pi_{Y}$-projections $\{\mathcal{Y}^{+}_{j}\}_{j}$ and from the knowledge of the contact form $\beta_{Y}^{\partial_{+}}=_{\mathsf{def}}\beta_{Y}||_{\partial_{1}^{+}Y(v_{Y})}$. Specifically, for an odd $j\in[1,2n-1]$,
@@
$\mathcal{Z}_{j}^{+}=_{\mathsf{def}}\partial(\mathcal{X}_{j}^{+})\setminus(\Psi\big{(}\partial_{j}^{+}X(v_{X})\big{)}\cup\mathcal{Y}_{j}^{+}).$
We claim that, for an odd $j$, $\int_{\mathcal{Z}_{j}^{+}}(d\beta_{Y})^{\frac{2n+1-j}{2}}=0$ since $v_{Y}\in\ker(d\beta_{Y})$ and $\mathcal{Z}_{j}^{+}$ of dimension $2n+1-j$ consists of portions of $v_{Y}$-trajectories.
@@
$=\int_{\partial^{+}_{j}X(v_{X})}\Psi^{\ast}\big{(}d\beta_{Y}\big{)}^{\frac{2n+1-j}{2}}=\int_{\Psi\big{(}\partial^{+}_{j}(X)(v_{X})\big{)}}\big{(}d\beta_{Y}\big{)}^{\frac{2n+1-j}{2}}.$
@@
This validates the claim of Theorem 6.3 for odd $j$.
@@
$=\int_{\partial_{j-1}^{+}X(v_{X})}\Psi^{\ast}\big{(}(d\beta_{Y})^{\frac{2n+2-j}{2}}\big{)}=\int_{\Psi\big{(}\partial^{+}_{j-1}(X)(v_{X})\big{)}}(d\beta_{Y})^{\frac{2n+2-j}{2}}.$

### v6

@@ 6. Non-squeezing of $2n$-symplectic volumes, Morse wrinkles of Reeb fields, and holography
Using that $\beta_{1}=\Psi^{\ast}(\beta_{2})$, we see that $\Psi_{\ast}(v_{1})=v_{2}|_{\Psi(X_{1})}$. Moreover, the combinatorial tangency pattern (see the discussion preceding formula (2.2)) of every $v_{1}$-trajectory $\gamma_{1}$ to $\partial_{1}X_{1}$ coincides with the tangency pattern to $\Psi(\partial_{1}X)$ of the portion $\Psi(\gamma_{1})$ of the unique $v_{2}$-trajectory $\gamma_{2}\supset\Psi(\gamma_{1})$. Note that the tangency pattern of $\gamma_{2}$ to $\Psi(\partial_{1}X)$ could be different.
@@
Now we are going to validate the equatorial inequality (see Fig. 1 for all the ingredients of the constructions below). Let $\mathcal{W}_{2}(\Psi)$ denote the $2n$-dimensional set comprised of $v_{2}$-trajectories through the locus $\Psi(\partial_{2}X_{1}(v_{1}))\coprod\partial_{2}X_{2}(v_{2})$. Consider the set (see Fig.  1)
@@
along which the continuous dependence of the $v_{2}$-flow on the initial conditions in $\Psi(\partial_{1}^{+}X_{1}(v_{1}))$ holds. Let $\mathcal{V}^{\circ}_{1,\kappa}(\Psi)$ be a connected component of the set $\mathcal{V}_{1}^{\circ}(\Psi)$, and let $\mathcal{V}_{1,\kappa}(\Psi)$ be the closure of $\mathcal{V}_{1,\kappa}^{\circ}(\Psi)$. We denote by $\partial^{+}_{1}X^{\circ}_{1,\kappa}(v_{1})\subset\partial_{1}^{+}X(v_{1})$ the $\Psi$-preimage of $\mathcal{V}^{\circ}_{1,\kappa}(\Psi)$.
@@
Let $\mathcal{Y}_{\kappa}(\Psi)=_{\mathsf{def}}\mathcal{X}_{\kappa}(\Psi)\cap\partial_{1}^{+}X_{2}(v_{2})$ be the set where the $(-v_{2})$-trajectories through $\mathcal{V}_{1,\kappa}(\Psi)$ terminate. Then the boundary of the set $\mathcal{X}_{\kappa}(\Psi)$ consists of three parts: $\mathcal{V}_{1,\kappa}(\Psi)),\,\hfill\penalty-10000\mathcal{Y}_{\kappa}(\Psi)$, and the rest, denoted by $\mathcal{Z}_{\kappa}(\Psi)$. The $2n$-dimensional Lebesgue measures of mutual intersections of these three parts is zero.
@@
Adding the inequalities for all $\kappa$’s and using that $\{\partial_{1}^{+}X^{\circ}_{1,\kappa}(v_{1})\}_{\kappa}$ form a partition of $\partial_{1}^{+}X_{1}(v_{1})$, modulo the set of $2n$-dimensional measure zero, we get
@@
    $c^{\bullet}(\Psi,v_{\beta_{2}})\cdot\int_{\partial_{2}Y(v_{Y})}d\beta_{Y}\wedge(d\beta_{Y})^{n-1},$
where $\Gamma(\sim)$ is the Gamma function. In particular, for $X$ being a ball of radius $r$, we get the following constraints on the injectivity radius $r$ of a contact embedding: πn+12Γ(n+32)  r^2n+1 ≤∫_Y β_Y ∧(dβ_Y)^n, r ≤diam_R(β_Y), πnΓ(n + 1)  r^2n  ≤  c^∙(Ψ, v_β_2)⋅∫_∂_2Y(v_Y) dβ_Y ∧(dβ_Y)^n-1. $\diamondsuit$
@@
Consider a solid $X$ shaped as a sand clock (see Fig. 2) which is vertically alined with the $z$-axis in the xyzxyz-space $\mathbb{R}^{3}$. Let $\beta=dz+x\,dy$. We consider a $2$-parameter family of transformations $\{A_{\lambda\mu}:x\to\lambda x,y\to\lambda y,z\to\mu z\}_{\lambda,\mu\in\mathbb{R}_{+}}$. We compare two contact $3$-folds $(X,\beta)$ and $(A_{\lambda\mu}(X),\beta)$. Then $c^{\bullet}(\Psi,v_{2})\geq 2$.
@@
For any boundary generic Reeb vector field $v_{\beta}$ on a compact $(2n+1)$-manifold $X$, we have $\partial_{j}X(v_{\beta})=\partial(\partial_{j-1}^{+}X(v_{\beta}))$ for all $j\geq 1$. Thus, for an odd $j$, by Stokes’ theorem,
$\int_{\partial_{j}X(v_{\beta})}(d\beta)^{\frac{2n+1-j}{2}}=\int_{\partial_{j-1}^{+}X(v_{\beta})}d\big((d\beta)^{\frac{2n+1-j}{2}}\big)=\int_{\partial_{j-1}^{+}X(v_{\beta})}\mathbf{0}\;=\;0.$
@@
$\kappa_{j}^{+}(\beta)$ $=_{\mathsf{def}}\;$ $\int_{\partial_{j}^{+}X(v_{\beta})}(d\beta)^{\frac{2n+1-j}{2}}\;\;\;(\equiv\kappa_{j+1}(\beta)),$
(6.5) $\kappa_{j}(\beta)$ = $=\;0$  $\text{\; for an odd }j\in[1,2n-1].$
@@
(6.6) $\kappa_{j}(\beta)$ $=_{\mathsf{def}}\;$ $\int_{\partial_{j}X(v)}\beta\wedge(d\beta)^{\frac{2n-j}{2}}\;\;\;(\equiv\kappa^{+}_{j-1}(\beta))\text{\; for an even }j\in[2,2n],$
Of course, these quantities depend on the choice of the contact form $\beta$ that represents a given contact structure $\xi$ and whose Reeb vector field $v_{\beta}$ is boundary generic. They measure how “wrinkled” $\partial_{1}X$ is in relation to the Reeb flow. By the very definition of $\kappa_{j}^{+}(\beta),\kappa_{j}(\beta)$, they belong to the category of “boundary data”.
With these quantities in place, note that the inequality (6.2) can be written as
$\kappa^{+}_{0}(\beta)\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\kappa^{+}_{1}(\beta),$
and the inequality (6.3) as
$\kappa^{+}_{0}(\beta)\leq\mathsf{diam}_{\mathcal{R}}(\beta)\cdot\kappa_{2}(\beta),$
where $\kappa^{+}_{0}(\beta)=_{\mathsf{def}}\int_{X}\beta\wedge(d\beta)^{n}.$
@@ 6.1. Invariance of the volumes $\{\kappa_{j}(\beta)\}$ under special deformations of $\beta$
Let $\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ be a family of contact forms on a compact connected smooth $(2n+1)$-manifold $X$ with boundary.
@@
for a $t$-family of smooth functions $\eta_{t}:\partial_{1}X\to\mathbb{R}$.
$\bullet$ We say the family $\mathcal{B}$ is variationally boundary (2n+1j)(2n+1-j)-exact if, for some smooth family of (2nj)(2n-j)-forms $\{\zeta_{t}\}_{t\in[0,1]}$ on $\partial_{1}X$,
(6.7) $\Big\{\frac{\partial}{\partial t}\Big[\beta_{t}\wedge(d\beta_{t})^{\frac{2n-j}{2}}\Big]\Big\}\Big|_{\partial_{1}X}=d\zeta_{t},\text{ where }j\in[2,2n]\text{ is even, }$
[…]
@@
Any variationally boundary exact family $\mathcal{B}$ of contact forms $\{\beta_{t}\}$ is variationally boundary (2n+1j)(2n+1-j)-exact for all $j$.
@@
Note that, for any $t$-family differential forms, their exterior differential $d=_{\mathsf{def}}d_{t}$ commutes with the partial derivative operator $\frac{\partial}{\partial t}$ and with the restriction of forms to $\partial_{1}X$.
Consider the derivative
@@
Since we proved that $\frac{\partial}{\partial t}\Big[(d\beta_{t})^{k}\Big]\Big|_{\partial_{1}X}=0$, the restriction of this formula to $\partial_{1}X$ collapses to $\frac{\partial}{\partial t}\Big[\beta_{t}\Big]\wedge(d\beta_{t})^{k}\Big|_{\partial_{1}X}$. Replacing $\frac{\partial}{\partial t}\Big[\beta_{t}\Big]\Big|_{\partial_{1}X}$ with $d\eta_{t}$ once more, we get
@@
Let $\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ be a variationally boundary (2n+1j)(2n+1-j)-exact family of contact forms on a compact connected smooth $(2n+1)$-manifold $X$. Assume that their Reeb vector fields $v_{\beta_{t}}$ are boundary generic for $t=0,1$, and for almost all $t\in(0,1)$.
$\bullet$ Then, for $j\equiv 0\mod 2$, the integrals $\kappa_{j}(\beta_{t})$ from (6.6) and $\kappa_{j-1}^{+}(\beta_{t})$ from (6) are $t$-independent. If, in addition, the property (6.9) holds, then $\kappa^{+}_{j}(\beta_{t})$ is also $t$-independent for all $j\equiv 0\mod 2$.
$\bullet$ If $\mathcal{B}=\{\beta_{t}\}_{t\in[0,1]}$ is variationally boundary exact family, then $\{\kappa_{j}(\beta_{t})\}$ are $t$-independent for all even $j\in[2,2n]$. If in addition, (6.9) holds for the function $\eta_{t}$, then $\{\kappa^{+}_{j}(\beta_{t})\}$ are $t$-independent for all $j\equiv 0\mod 2$.
@@
$\int_{\mathcal{Y}_{j}}(d_{t}\beta_{t})^{s+1}\;+\int_{\mathcal{Y}_{j}}\frac{\partial}{\partial t}\Big[\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big]\wedge dt$
Each of these two integrals vanishes. Indeed, $dt$ is not present in the $(2s+2)$-form $(d_{t}\beta_{t})^{s+1}$. Thus its restriction to the $(2s+2)$-chain $\mathcal{Y}_{j}\subset\partial_{1}X\times[0,1]$ vanishes. The second integral is also zero by the variationally boundary (2n+1j)(2n+1-j)-exact hypothesis:
$\int_{\mathcal{Y}_{j}}\frac{\partial}{\partial t}\big(\beta_{t}\wedge(d_{t}\beta_{t})^{s}\big)\wedge dt\stackrel{{\scriptstyle Fubini}}{{=}}\int_{[0,1]}\Big(\int_{\partial_{j}X(v_{\beta_{t}})}\frac{\partial}{\partial t}\Big[\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big]\Big)dt$
$\int_{[0,1]}\Big(\int_{\partial_{j}X(v_{\beta_{t}})}d_{t}\zeta_{t}\Big)dt=0,$
@@
Using Stokes’ theorem identity $\{\kappa_{j}(\beta)=\kappa_{j-1}^{+}(\beta)\}$ ($j\equiv 0\mod 2$), we conclude that, for odd $j\in[1,2n-1]$, the quantities $\{\kappa_{j}^{+}(\beta)\}_{j}$ are $\mathcal{B}$-deformation invariants as well.
If we add property (6.9) to property (6.7), then a similar computation shows that $\kappa^{+}_{j}(\beta_{1})=\kappa^{+}_{j}(\beta_{0})$ for an even $j$. This time, Stokes’ theorem is applied to $\mathcal{Y}_{j}^{+}$. Its boundary consists of tree parts: $\partial_{j}^{+}X(v_{0}),\partial_{j}^{+}X(v_{1})$, and their complement $\delta\mathcal{Y}_{j}^{+}=\mathcal{Y}_{j+1}$. Therefore,
@@
$\int_{\mathcal{Y}^{+}_{j}}(d_{t}\beta_{t})^{s+1}\;+\int_{\mathcal{Y}^{+}_{j}}\frac{\partial}{\partial t}\Big[\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big]\wedge dt\;\pm\;\int_{\delta\mathcal{Y}^{+}_{j}}\beta_{t}\wedge(d_{t}\beta_{t})^{s}.$
@@
$\kappa^{+}_{j}(\beta_{1})-\kappa^{+}_{j}(\beta_{0})=\int_{[0,1]}\Big(\int_{\partial^{+}_{j}X(v_{\beta_{t}})}\frac{\partial}{\partial t}\Big[\beta_{t}\wedge(d_{t}\beta_{t})^{s}\Big]\Big)dt$
$\int_{[0,1]}\Big(\int_{\partial^{+}_{j}X(v_{\beta_{t}})}d_{t}\zeta_{t}\Big)dt\;\stackrel{{\scriptstyle Stokes}}{{=}}\;\int_{[0,1]}\Big(\int_{\partial_{j+1}X(v_{\beta_{t}})}\zeta_{t}\Big)dt=0$
@@
(6.10) $\mathcal{K}_{j}^{+}(\xi)$ $=_{\mathsf{def}}\;\inf_{\{\beta\leadsto\xi\;|\;\mathsf{av}_{\mathcal{R}}(\beta)=1\}}\big\{|\kappa_{j}^{+}(\beta)|\big\}$  $\text{\; for an odd }j\in[1,2n-1],$
$\mathcal{K}_{j}^{+}(\xi)$ $=_{\mathsf{def}}\;\inf_{\{\beta\leadsto\xi\;|\;\mathsf{av}_{\mathcal{R}}(\beta)=1\}}\big\{|\kappa_{j}^{+}(\beta)|\big\}$  $\text{\; for an even }j\in[2,2n],$
(6.11) $\mathcal{K}_{j}(\xi)$ $=_{\mathsf{def}}\;\inf_{\{\beta\leadsto\xi\;|\;\mathsf{av}_{\mathcal{R}}(\beta)=1\}}\big\{|\kappa_{j}(\beta)|\big\}$  $\text{\; for an even }j\in[2,2n].$
@@
Figure 2. By Theorem 6.3, $\kappa_{1}^{+}(\beta_{X})=\kappa_{2}(\beta_{X})=-\int_{\mathcal{Y}_{1}^{+}}d\beta_{Y}$. The figure illustrates the geometry of $3$-folds $X$ and $Y$ and of an embedding $\Psi:X\to Y$ that leads to this formula. Although the relevant for the formula loci are: the surface $\Psi(\partial_{1}^{+}X)$, the solid $\mathcal{X}^{+}_{1}\subset Y$, and the surface $\mathcal{Y}^{+}_{1}\subset\partial_{1}^{+}Y$, they are not shown in the figure due to artistic limitations; instead, we have chosen to show similar loci of one dimension lower: the curve $\Psi(\partial_{2}^{+}X)$, the surface $\mathcal{X}^{+}_{2}$, and the curve $\mathcal{Y}^{+}_{2}$.
@@
Then the integrals $\{\kappa_{j}^{+}(\beta_{X})\}_{j}$ from (6) and $\{\kappa_{j}(\beta_{X})\}_{j}$ from (6.6) can be recovered from the $\Pi_{Y}$-projections $\{\mathcal{Y}^{+}_{j}\}_{j}$ and from the knowledge of the contact form $(\beta_{Y})^{\partial_{+}}=_{\mathsf{def}}\beta_{Y}||_{\partial_{1}^{+}Y(v_{Y})}$. Specifically, for an odd $j\in[1,2n-1]$,
@@
$\mathcal{Z}_{j}^{+}=_{\mathsf{def}}\partial(\mathcal{X}_{j}^{+})\setminus\big(\Psi\big(\partial_{j}^{+}X(v_{X})\big)\cup\mathcal{Y}_{j}^{+}\big).$
We claim that, for an odd $j$, $\int_{\mathcal{Z}_{j}^{+}}(d\beta_{Y})^{\frac{2n+1-j}{2}}=0$ since $v_{Y}\in\ker(d\beta_{Y})$ and $\mathcal{Z}_{j}^{+}$ of dimension 2n+1j2n+1-j consists of portions of $v_{Y}$-trajectories.
@@
$\int_{\partial^{+}_{j}X(v_{X})}\Psi^{\ast}\big(d\beta_{Y}\big)^{\frac{2n+1-j}{2}}=\int_{\Psi\big(\partial^{+}_{j}(X)(v_{X})\big)}\big(d\beta_{Y}\big)^{\frac{2n+1-j}{2}}.$
@@
This validates the claim of Theorem 6.3 for an odd $j$.
@@
$\int_{\partial_{j-1}^{+}X(v_{X})}\Psi^{\ast}\big((d\beta_{Y})^{\frac{2n+2-j}{2}}\big)=\int_{\Psi\big(\partial^{+}_{j-1}(X)(v_{X})\big)}(d\beta_{Y})^{\frac{2n+2-j}{2}}.$